OpenStudy (anonymous):

What is the solution of the proportion... w + 14/ 4w + 6 = 3/4

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Since you described it as a proportion, I am guessing that this is what you meant: \(\dfrac{w+14}{4w+6}=\dfrac{3}{4}\)

OpenStudy (anonymous):

yes thats what I meant

OpenStudy (anonymous):

The equation editor can help you to avoid confusion. So, to solve a proportion, you cross multiply.

OpenStudy (anonymous):

i know how to cross multiply, but im confused with the w and the 4w

OpenStudy (anonymous):

What did you get when you cross multiplied?

OpenStudy (anonymous):

18 and 56?

OpenStudy (anonymous):

 I wish they would allow the use of colors. Instead I will demonstrate in a few steps....

OpenStudy (anonymous):



OpenStudy (anonymous):

Then distribute.

OpenStudy (anonymous):

so then you get.. 4w + 56 = 12w +18

OpenStudy (anonymous):

Yes, and then solve for w.

OpenStudy (anonymous):

So the answer is: w=4.5 thank you

OpenStudy (anonymous):

REcheck your calculations....

OpenStudy (anonymous):

I got w=4.75

OpenStudy (anonymous):

oh i found my mistake i got 36/8 instead of 38/8

OpenStudy (anonymous):

Great Job! You can always check your own answers by plugging your answer for x back into the original equation. \[\dfrac{4.75+14}{4w+6}=\dfrac{3}{4}\] \[\dfrac{4.75+14}{4(4.75)+6}=\dfrac{3}{4}\] \[\dfrac{18.75}{19+6}=\dfrac{3}{4}\] \[\dfrac{18.75}{25}=\dfrac{3}{4}\] \[0.75=0.75\]